OpenStudy (anonymous):

PLease help If f(x) = |(x2 − 9)(x2 + 1)|, how many numbers in the interval [–1, 1] satisfy the conclusion of the Mean Value Theorem?

OpenStudy (anonymous):

I understand mean value theorem but I dont know where to start

zepdrix (zepdrix):

You understand mean value theorem? Ok great. So what slope value does the `secant line` that connects the ends points give us?

zepdrix (zepdrix):

\(\large\rm \frac{f(1)-f(-1)}{1-(-1)}=?\)

OpenStudy (anonymous):

solving that would give me f'(c)

OpenStudy (anonymous):

Im just confused how absolute value effects the entire thing

zepdrix (zepdrix):

Yes, but give me the number first for the slope, we'll deal with derivative later.

OpenStudy (anonymous):

(|-8/2|-0)/(2)

zepdrix (zepdrix):

Woops, shouldn't get 0 in the numerator. I think you forgot to square your -1, yes?

OpenStudy (anonymous):

i did haha

OpenStudy (anonymous):

(|-16|-|-16|)/2

zepdrix (zepdrix):

Ok good good good. So we're getting \(\large\rm \frac{f(1)-f(-1)}{1-(-1)}=0\)

zepdrix (zepdrix):

So notice, this problem boils down to a special case of the Mean Value Theorem, known as Rolle's Theorem. We're looking for zero slopes in this interval, better known as `critical points`.

zepdrix (zepdrix):

So now we at least know what we're looking for :) Derivative is going to be a little bit of a pain. Let me remind you of absolute value derivative.

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}|x|=\frac{x}{|x|}\]If our inner function is more than just x, don't forget to chain rule,\[\large\rm \frac{d}{dx}|stuff|=\frac{stuff}{|stuff|}\cdot(stuff)'\]

OpenStudy (anonymous):

so i have to get f"(x)?

zepdrix (zepdrix):

So it's going to look pretty scary, should simply down nicely though.\[\large\rm \frac{d}{dx}|(x^2-9)(x^2+1)|=\frac{(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot[(x^2-9)(x^2+1)]'\]

zepdrix (zepdrix):

No... why would you need f''?

zepdrix (zepdrix):

We're looking for critical points :o We don't care about inflection or anything like that.

OpenStudy (anonymous):

OKay sorry

zepdrix (zepdrix):

I know it looks like a big mess, but hopefully you can at least see how it follows the rule that I posted for absolute derivative.

zepdrix (zepdrix):

So now we have to multiply by the derivative of the inner function,\[\rm \frac{d}{dx}|(x^2-9)(x^2+1)|=\frac{(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot\color{royalblue}{[(x^2-9)(x^2+1)]'}\]We'll have to product rule here.

OpenStudy (anonymous):

solving it now

OpenStudy (anonymous):

(4 x (-9+x^2) (-4+x^2) (1+x^2))/(abs((-9+x^2) (1+x^2)))

OpenStudy (anonymous):

looks even messier lol

zepdrix (zepdrix):

I can't quite figure out what you did :( hmm

OpenStudy (anonymous):

i got the derivative 4x(x^2-4)

OpenStudy (anonymous):

I then multiplied it by the main function

zepdrix (zepdrix):

\[\rm =\frac{(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot\color{royalblue}{[(x^2-9)(x^2+1)]'}\]Product rule:\[\rm =\frac{(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot\color{royalblue}{[(x^2-9)'(x^2+1)+(x^2-9)(x^2+1)']}\]That's the setup yes?

zepdrix (zepdrix):

Oh oh oh, don't multiply stuff together :( yikes

OpenStudy (anonymous):

yes but i thought you got the derivative and then multiplied

zepdrix (zepdrix):

Finishing up product rule gives us,\[\rm =\frac{(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot\color{royalblue}{[(2x)(x^2+1)+(x^2-9)(2x)]}\]Factor a (2x) out of each term,\[\large\rm =\frac{(2x)(x^2-9)(x^2+1)}{|(x^2-9)(x^2+1)|}\cdot\left[x^2+1+x^2-9\right]\]Oh oh oh nevermind :) I see where your 4 is coming from. Ok ok good good good

zepdrix (zepdrix):

\[\large\rm =\frac{(4x)(x^2-9)(x^2+1)(x^2-4)}{|(x^2-9)(x^2+1)|}\]

zepdrix (zepdrix):

This is our derivative function. We're looking for zero-slope,\[\large\rm 0=\frac{(4x)(x^2-9)(x^2+1)(x^2-4)}{|(x^2-9)(x^2+1)|}\]Which leads to the numerator equaling zero,\[\large\rm 0=4x(x^2-9)(x^2+1)(x^2-4)\]

zepdrix (zepdrix):

Solve for x, and then determine which of those critical points you're allowed to use. (Only the ones within the interval [-1,1].)

zepdrix (zepdrix):

The derivative says "how many", they didn't actually ask you to find them, which leads me to believe that there is probably an easier way to do this problem XD lol

zepdrix (zepdrix):

the directions say*

OpenStudy (anonymous):

haha. The work was killer but Im glad we did it that way

OpenStudy (anonymous):

is it only 1 root then? x=0?

zepdrix (zepdrix):

yay good job \c:/ x=3,-3 and x=2,-2 and x=0 only 0 lies in the interval

OpenStudy (anonymous):

Youre the best. Thanks!

OpenStudy (anonymous):

not to butt in but i think you are doing way way too much work here

OpenStudy (anonymous):

:O. Can you tell me the shorter way haha

OpenStudy (anonymous):

this is your function right?\[f(x)= |(x^2 − 9)(x^2 + 1)|\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the interval is \([-1,1]\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

on that interval, \(x^2-9\) is negative so \(|x^2-9|=9-x^2\) as far as you are concerned

OpenStudy (anonymous):

and of course \(x^2+1\geq 0\) for all \(x\) so your function is just \[(9-x^2)(x^2+1)\] which is a lot easier to deal with

zepdrix (zepdrix):

True, heh

OpenStudy (anonymous):

haha wish i knew that. Thanks for pointing it out